OpenStudy (anonymous):

Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

ok so the formula for the curcimfence is 2*pi*r for the area is pi*r^2

OpenStudy (alexandervonhumboldt2):

circle A: circumference: 25.12 inches. thus 2*4*pi=25.12 pi=3.14 area: pi*4^2=50.24 pi=50.24/16=3.14

OpenStudy (alexandervonhumboldt2):

circle B: circumference: 2*1.5*pi=9.42 pi=9.42/3=3.14 area: 1.5^2*pi=7.065 pi=3.14

OpenStudy (alexandervonhumboldt2):

The number Pi is always the same. In any circle, pi is equal to 22/7 which is approximately 3.14

OpenStudy (alexandervonhumboldt2):

did you understood @geny55 ?

OpenStudy (anonymous):

yes @AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

so whats the answer for part A? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

circle A: circumference: 25.12 inches. thus 2*4*pi=25.12 pi=3.14 circumference: 2*1.5*pi=9.42 pi=9.42/3=3.14

OpenStudy (anonymous):

does circle A mean Part A?

OpenStudy (alexandervonhumboldt2):

part b: area: pi*4^2=50.24 pi=50.24/16=3.14 area: 1.5^2*pi=7.065 pi=3.14

OpenStudy (anonymous):

wait so part A is : circumference: 25.12 inches. thus 2*4*pi=25.12 pi=3.14 circumference: 2*1.5*pi=9.42 pi=9.42/3=3.14 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

ok .. part c?

OpenStudy (alexandervonhumboldt2):

The number Pi is always the same. In any circle, pi is equal to 22/7 which is approximately 3.14

OpenStudy (anonymous):

ok thanks :)

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (anonymous):

are all of these answers in your own words?

OpenStudy (alexandervonhumboldt2):

of cousre!